# A Bayesian statistical analysis of 897 pulsar flux density spectra

### v1

@@ I Introduction
Pulsars, the highly-magnetized and rapidly rotating remnants of massive stellar explosions, have been a source of fascination since their serendipitous discovery in 1967 (hewish_observation_1968). These compact neutron stars are enveloped by magnetospheres filled with electron-positron plasma that co-rotates with the star (goldreich_pulsar_1969). From our vantage point on Earth, they are observed as precise, periodic pulses of radio emission, akin to a cosmic lighthouse, each time their beamed radiation sweeps across our line of sight. This population of celestial clocks includes both isolated pulsars and those in binary systems (kramer_double_2008), serving as natural laboratories for studying extreme physics. Despite decades of study, the fundamental physical mechanisms governing their emission remain elusive, though they are thought to be intimately linked to properties such as rotation, magnetic field configuration, internal structure, and the surrounding environment.
@@
The quest to understand the pulsar emission mechanism has been pursued through a variety of methodological approaches over the past several decades. These efforts can be broadly categorized: some studies focus on in-depth analysis of remarkable individual objects to uncover specific physical phenomena (karastergiou_transient_2011; desvignes_radio_2019), while others seek to generalize findings by examining populations of pulsars that share common properties, such as age or magnetic field strength (johnston_galactic_2020; lower_impact_2021). A third, crucial approach involves large-scale statistical studies that aim to identify macroscopic trends and correlations among observable parameters across hundreds of pulsars (e.g., sieber_pulsar_1973; Jankowski_2018).
@@
A critical avenue for understanding pulsar emission is the study of their integrated radio flux density spectra—the variation of flux density with observing frequency. Measuring the flux density not only allows for the calculation of radio luminosity and energy output but also provides key insights into the underlying emission processes. However, obtaining accurate measurements is challenging. Interstellar scintillation and instrumental systematic errors can cause flux density to fluctuate by factors of 2 up to an order of magnitude over timescales ranging from weeks to months (swainston_discovery_2021). Consequently, high-precision measurements require observations that extend beyond the scintillation timescale, coupled with rigorous error analysis and advancing calibration techniques. Thanks to such sustained efforts, flux density measurements for thousands of pulsars have now been accumulated in various catalogues and databases (e.g., manchester_australia_2005; swainston_pulsar_spectra_2022).
@@
It has long been thought that pulsar spectra often approximately follow a power-law distribution, with a characteristic spectral index $\alpha$ ranging from 2-2 to 1-1 (e.g., Lorimer_1995; xilouris_emission_1996; maron_pulsar_2000; bates_pulsar_2013; Jankowski_2018). However, a significant number of pulsars exhibit more complex spectral shapes, including breaks or turn-overs (sieber_pulsar_1973). The standard analytical models used to describe these spectra include, beyond the simple power law, the broken power law and the log-parabolic spectrum. The latter two are particularly useful for modeling gigahertz-peaked spectra (GPS), where the flux density peaks around 1 GHz (xilouris_emission_1996; bates_pulsar_2013; Dembska_2014).
@@
Previous population studies of pulsar flux density spectra have laid important groundwork. For instance, Jankowski_2018 performed a spectral analysis of 441 pulsars, concluding that the simple power law was dominant (79%) based on the Akaike Information Criterion (AIC) for model comparison. More recently, Posselt_2023 presented flux measurements to 1,170 pulsars and performed power-law fits of pulsar spectra. We build upon previous work in the following ways. First, from an expanded dataset of 3,345 pulsars, we select 897 high-quality sources with at least four flux density measurements spanning a factor of two in frequency. Second, we employ a Bayesian framework for model comparison, in addition to replicating the frequentist approach of Jankowski_2018. Third, we systematically evaluate six spectral models: the simple power law, broken power law, low-frequency turn-over power law, high-frequency cut-off power law, double turn-over spectrum, and log-parabolic spectrum.
@@ II.1 Data compilation
Our primary data source is the pulsar_spectra [^1] https://github.com/NickSwainston/pulsar_spectra Python package (version 2.1.0) swainston_pulsar_spectra_2022, a well-maintained and comprehensive catalog of pulsar flux densities. We manually reviewed the literature compiled within this repository and removed entries originating from uncalibrated publications. To incorporate recent measurements and expand this base dataset, we supplemented it with data from the following key studies: sieber_pulsar_1973; maron_pulsar_2000; levin_radio-loud_2010; anderson_multi-wavelength_2012; gordon_quick_2021; Spiewak_2022; anumarlapudi_characterizing_2023; Posselt_2023. A complete list of all references included in our dataset is provided in the spectral_fit [^2] https://github.com/QZGao/spectral_fit repository.
@@
For model fitting and comparison, we apply the quality criteria adopted by Jankowski_2018. We select only those pulsars with at least four distinct frequency measurements spanning a frequency range of at least a factor of two. This result in a final high-quality sample of 897 pulsars that meet these requirements. Notably, this sample is more than twice the size of the 441 pulsars analyzed in the foundational study of Jankowski_2018, providing a substantially larger basis for robust statistical analysis.
@@ II.2 Bayesian model selection
We apply Bayesian inference to fit the flux density spectra of each pulsar in our data set. Our analysis follows the methodology described in youDeterminationBirthmassFunction2025. To compute posterior distributions for the model parameters and the Bayesian evidence for each model, we employ the dynamic nested sampling algorithm implemented in the Python package dynesty.
Figure 1: Distributions of relative uncertainties in flux density measurements from several major references and of the full 2,310-pulsar dataset. In Hobbs_2004a, 38.2% of the measurements adopt an assumed 10% relative uncertainty where no estimates were reported. Jankowski_2018 derive uncertainties using modulation indices, producing a broad distribution. Johnston_2018 do not provide uncertainties, so a 20% relative uncertainty was assigned by pulsar_spectra. Posselt_2023 employ an unconventional flux density measurement method that yields relatively small uncertainties for most pulsars. Across all literature, 6.3% of the measurements have relative uncertainties of $\sim$10%, 3.5% have $\sim$20%, and 13.2% have $\sim$50%.
A key challenge in spectral modeling is the variability and reliability of reported flux density uncertainties. Uncertainty estimates differ widely across studies: some authors adopt fixed values—such as 50%, 20%, or 10%—either alone or added in quadrature to measurement errors; others average measurements obtained over multiple epochs and use the standard deviation; and more recent work sometimes adopts the modulation index. These methodological differences reflect varying instrumentation, observational strategies, and calibration procedures, as discussed in Gitika_2023. The absence of a unified standard is further illustrated in Fig. 1, which shows the distribution of relative uncertainties across selected studies and the full dataset.
In many publications, it is unclear whether the reported uncertainties adequately account for temporal variability or systematic effects. Underestimated uncertainties are particularly problematic for spectral fitting, as their quadratic contribution to the likelihood can disproportionately influence model selection. To mitigate this, Jankowski_2018 adopted the Huber loss function to downweight outliers. However, this frequentist approach is not directly compatible with our Bayesian framework.
@@
The upper bound on $e_{\text{quad}}$ is motivated by the distribution of relative errors shown in Fig. 1.
@@ II.3 Spectral models
To characterize the diverse spectral shapes present in our pulsar sample, we evaluate six phenomenological models. These include the five models considered in Jankowski_2018 and the double turn-over model introduced by swainston_pulsar_spectra_2022. The models are summarized below.
@@
Originally proposed to account for spectral breaks (sieber_pulsar_1973), the broken power law model consists of two power-law segments joined at a break frequency $\nu_{b}$:
$S(\nu)=\begin{cases}b\nu^{\alpha_{1}},&\textrm{if }\nu\leq\nu_{b},\\ b\nu^{\alpha_{2}}\nu_{b}^{\alpha_{1}-\alpha_{2}},&\textrm{otherwise,}\end{cases}$ (4)
where $\alpha_{1}$ and $\alpha_{2}$ are the spectral indices below and above the break.
The log-parabolic spectrum, used to identify spectral curvature and turn-overs (Kuzmin_2001; kijak_spectrum_2011), is written as
$S(\nu)=\exp\left(a\nu^{2}+b\nu+c\right),$ (5)
@@
Two additional models are derived from a general form that includes an absorption or cut-off term (sieber_pulsar_1973; malofeev_mean_1980):
$S(\nu)=b\nu^{\alpha}e^{-\tau(\nu)}\,.$ (6)
The low-frequency turn-over power law (Jankowski_2018; rajwade_gigahertz_2016) models the optical depth to produce a low-frequency turn-over, often attributed to thermal free–free absorption (kijak_spectrum_2011; wilson_tools_2013):
$S(\nu)=b\nu^{\alpha}\exp\left(\frac{\alpha}{\beta}\nu_{c}^{-\beta}\right),$ (7)
where $\nu_{c}$ is the turn-over frequency and $\beta\in[0,2.1]$ controls the smoothness of the transition.
The high-frequency cut-off power law incorporates a high-frequency cut-off by adopting an optical-depth term that produces a linear attenuation with frequency (kontorovich_high-frequency_2013). In this case, the optical depth takes the form $\tau(\nu)=-\ln\!\left(1-\nu/\nu_{c}\right)$, leading to
$S(\nu)=b\nu^{\alpha}\left(1-\frac{\nu}{\nu_{c}}\right),$ (8)
where $\nu_{c}$ is the cut-off frequency. Such high-frequency behaviour is generally interpreted as arising from the pulsar’s interaction with its environment rather than the intrinsic emission mechanism (kijak_low-frequency_2021; ardavan_radio_2024).
Motivated by earlier attempts to combine low- and high-frequency features (sieber_pulsar_1973), the double turn-over spectrum (swainston_pulsar_spectra_2022) includes both a low-frequency turn-over and a high-frequency cut-off:
$S(\nu)=b\nu^{\alpha}\left(1-\frac{\nu}{\nu_{c_{1}}}\right)\exp\left(\frac{\alpha}{\beta}\nu_{c_{2}}^{-\beta}\right),$ (9)
where $\nu_{c_{1}}$ and $\nu_{c_{2}}$ are the cut-off and turn-over frequencies, respectively.
To mitigate parameter covariance and reduce variance in the amplitude $b$, we follow established practice (swainston_pulsar_spectra_2022; Bilous_2016) by scaling frequency with a reference value $\nu_{0}$, defined as the geometric mean of the minimum and maximum observed frequencies for each pulsar. This transforms the frequency variable to a dimensionless ratio $x=\nu/\nu_{0}$. The prior distributions for all model parameters are listed in Table. 1.
@@
Low-frequency turn-over
power law
$\nu_{c}$ [MHz] $LU(\nu_{\min},\nu_{\max})$
@@
High-frequency cut-off
power law
$\nu_{c}$ [MHz] $LU(\nu_{\min},1.5\,\nu_{\max})$
@@ III Results
Figure 2: Spectrum of PSR J1809–1917, which was previously classified as a simple power law Jankowski_2018 (lower panel), but our Bayesian analysis identifies the low-frequency turn-over power law as the best model (upper panel), with a $\ln\mathrm{BF}$ of 150 in its favor. The solid orange line and the shaded region display the posterior predictive distribution and its associated uncertainty, with the dark and light colored region showing the 1-$\sigma$ and 3-$\sigma$ credibility interval, respectively. Solid error bars denote the uncertainties reported in the literature, while dashed error bars represent the rescaled uncertainties with inferred $e_{\text{fac}}$ parameters.
@@ III.1 Bayesian model selection
Best-fit model
$\ln\mathrm{BF}\geq 5$
(vs. second-best model)
$\ln\mathrm{BF}\geq 5$
(vs. simple power law)
@@
Solely relying on log evidence for pulsar spectral classification can be sometimes misleading when the evidence difference between models is small. We compute the log Bayes factor between each curved or broken model and the simple power law. A log Bayes factor of at least 5 is commonly interpreted as decisive support. As shown in Table. 3, 68.8% of pulsars are decisively better fit ($\ln\mathrm{BF}\geq 5$) by a curved or broken model than by a simple power law, indicating that such spectra are common. Even under a more stringent threshold ($\ln\mathrm{BF}\geq 8$), 65.5% of pulsars still favor one of the curved or broken models against the simple power law. Conversely, only 3 of the 121 pulsars that are best fit with a simple power law spectrum has a log Bayes factor over 5 against other spectral models. This suggests that, even when the simple power law yields the highest evidence, its preference is not very strong.
We repeat the analysis using only pulsars with at least ten flux density measurements. With this larger number of data points, the spectral shape is more tightly constrained. Under this stricter requirement, 76.4% (584 out of 764) of categorized pulsars decisively favor ($\ln\text{BF}\geq 5$) curved or broken models over the simple power law, indicating that the prevalence of such spectra is not affected by the small number of measurements. Therefore, we conclude that curved or broken spectra are significantly more prevalent than reported by Jankowski_2018.

### v2

# Beyond the Simple Power Law: A Bayesian Analysis of 897 Pulsar Spectra

@@ I Introduction
Pulsars, the highly-magnetized and rapidly rotating remnants of massive stellar explosions, have been a source of fascination since their serendipitous discovery in 1967 (Hewish et al., 1968). These compact neutron stars are enveloped by magnetospheres filled with electron-positron plasma that co-rotates with the star (Goldreich and Julian, 1969). From our vantage point on Earth, they are observed as precise, periodic pulses of radio emission, akin to a cosmic lighthouse, each time their beamed radiation sweeps across our line of sight. This population of celestial clocks includes both isolated pulsars and those in binary systems (Kramer and Stairs, 2008), serving as natural laboratories for studying extreme physics. Despite decades of study, the fundamental physical mechanisms governing their emission remain elusive, though they are thought to be intimately linked to properties such as rotation, magnetic field configuration, internal structure, and the surrounding environment.
@@
The quest to understand the pulsar emission mechanism has been pursued through a variety of methodological approaches over the past several decades. These efforts can be broadly categorized: some studies focus on in-depth analysis of remarkable individual objects to uncover specific physical phenomena (Karastergiou et al., 2011; Desvignes et al., 2019), while others seek to generalize findings by examining populations of pulsars that share common properties, such as age or magnetic field strength (Johnston et al., 2020; Lower et al., 2021). A third, crucial approach involves large-scale statistical studies that aim to identify macroscopic trends and correlations among observable parameters across hundreds of pulsars (e.g., Sieber, 1973; Jankowski et al., 2018).
@@
A critical avenue for understanding pulsar emission is the study of their integrated radio flux density spectra—the variation of flux density with observing frequency. Measuring the flux density not only allows for the calculation of radio luminosity and energy output but also provides key insights into the underlying emission processes. However, obtaining accurate measurements is challenging. Interstellar scintillation and instrumental systematic errors can cause flux density to fluctuate by factors of 2 up to an order of magnitude over timescales ranging from weeks to months (Swainston et al., 2021). Consequently, high-precision measurements require observations that extend beyond the scintillation timescale, coupled with rigorous error analysis and advancing calibration techniques. Thanks to such sustained efforts, flux density measurements for thousands of pulsars have now been accumulated in various catalogues and databases (e.g., Manchester et al., 2005; Swainston et al., 2022).
@@
It has long been thought that pulsar spectra often approximately follow a power-law distribution, with a characteristic spectral index $\alpha$ ranging from 2-2 to 1-1 (e.g., Lorimer et al., 1995; Xilouris et al., 1996; Maron et al., 2000; Bates et al., 2013; Jankowski et al., 2018). However, a significant number of pulsars exhibit more complex spectral shapes, including breaks or turn-overs (Sieber, 1973). The standard analytical models used to describe these spectra include, beyond the simple power law, the broken power law and the log-parabolic spectrum. The latter two are particularly useful for modeling gigahertz-peaked spectra (GPS), where the flux density peaks around 1 GHz (Xilouris et al., 1996; Bates et al., 2013; Dembska et al., 2014).
@@
Previous population studies of pulsar flux density spectra have laid important groundwork. For instance, Jankowski et al. (2018) performed a spectral analysis of 441 pulsars, concluding that the simple power law was dominant (79%) based on the Akaike Information Criterion (AIC) for model comparison. More recently, Posselt et al. (2023) presented flux measurements to 1,170 pulsars and performed power-law fits of pulsar spectra. We build upon previous work in the following ways. First, from an expanded dataset of 3,345 pulsars, we select 897 high-quality sources with at least four flux density measurements spanning a factor of two in frequency. Second, we employ a Bayesian framework for model comparison, in addition to replicating the frequentist approach of Jankowski et al. (2018). Third, we systematically evaluate six spectral models: the simple power law, broken power law, low-frequency turn-over power law, high-frequency cut-off power law, double turn-over spectrum, and log-parabolic spectrum.
@@ II.1 Data compilation
Our primary data source is the pulsar_spectra [^1] https://github.com/NickSwainston/pulsar_spectra Python package (version 2.1.0) Swainston et al. (2022), a well-maintained and comprehensive catalog of pulsar flux densities. We manually reviewed the literature compiled within this repository and removed entries originating from uncalibrated publications. To incorporate recent measurements and expand this base dataset, we supplemented it with data from the following key studies: Sieber (1973); Maron et al. (2000); Levin et al. (2010); Anderson et al. (2012); Gordon et al. (2021); Spiewak et al. (2022); Anumarlapudi et al. (2023); Posselt et al. (2023). A complete list of all references included in our dataset is provided in the spectral_fit [^2] https://github.com/QZGao/spectral_fit repository.
@@
For model fitting and comparison, we apply the quality criteria adopted by Jankowski et al. (2018). We select only those pulsars with at least four distinct frequency measurements spanning a frequency range of at least a factor of two. This result in a final high-quality sample of 897 pulsars that meet these requirements. Notably, this sample is more than twice the size of the 441 pulsars analyzed in the foundational study of Jankowski et al. (2018), providing a substantially larger basis for robust statistical analysis.
@@ II.2 Bayesian model selection
We apply Bayesian inference to fit the flux density spectra of each pulsar in our data set. Our analysis follows the methodology described in You et al. (2025). To compute posterior distributions for the model parameters and the Bayesian evidence for each model, we employ the dynamic nested sampling algorithm implemented in the Python package dynesty. While the evidence includes an Occam penalty for model complexity, we treat the highest-evidence model as a convenient classification label rather than a definitive choice. We assess the strength of statistical preference using Bayes factors.
Figure 1: Distributions of relative uncertainties in flux density measurements from several major references and of the full 2,310-pulsar dataset. In Hobbs et al. (2004), 38.2% of the measurements adopt an assumed 10% relative uncertainty where no estimates were reported. Jankowski et al. (2018) derive uncertainties using modulation indices, producing a broad distribution. Johnston and Kerr (2018) do not provide uncertainties, so a 20% relative uncertainty was assigned by pulsar_spectra (Swainston et al., 2022). Posselt et al. (2023) employ an unconventional flux density measurement method that yields relatively small uncertainties for most pulsars.
A key challenge in spectral modeling is the variability and reliability of reported flux density uncertainties. Uncertainty estimates differ widely across studies: some authors adopt fixed values—such as 50%, 20%, or 10%—either alone or added in quadrature to measurement errors; others average measurements obtained over multiple epochs and use the standard deviation; and more recent work sometimes adopts the modulation index. These methodological differences reflect varying instrumentation, observational strategies, and calibration procedures, as discussed in Gitika et al. (2023). The absence of a unified standard is further illustrated in Fig. 1, which shows the distribution of relative uncertainties across selected studies and the full dataset. For measurements that do not report accurate uncertainties, we adopt the assumed relative uncertainties suggested in the original papers (e.g., 10% from Hobbs et al. (2004); Frail et al. (2016)), rather than imposing new values. If no such estimates are provided, we assume a 50% relative uncertainty, a value widely adopted in earlier works (Jankowski et al., 2018; Sieber, 1973).
In many publications, it is unclear whether the reported uncertainties adequately account for temporal variability or systematic effects. Underestimated uncertainties are particularly problematic for spectral fitting, as their quadratic contribution to the likelihood can disproportionately influence model selection. To mitigate this, Jankowski et al. (2018) adopted the Huber loss function to downweight outliers. However, this frequentist approach is not directly compatible with our Bayesian framework.
@@
The upper bound on $e_{\text{quad}}$ is motivated by the distribution of relative errors shown in Fig. 1. Specifically, 96.5% of all reported relative uncertainties in Fig. 1 are $\leq 50\%$, motivating the 50% systematic term as a conservative limit.
@@ II.3 Spectral models
To characterize the diverse spectral shapes present in our pulsar sample, we evaluate six phenomenological models. These include the five models considered in Jankowski et al. (2018) and the double turn-over model introduced by Swainston et al. (2022). To mitigate parameter covariance, we scale frequency with a reference value $\nu_{0}$, defined as the geometric mean of the minimum and maximum observed frequencies. This transforms the frequency variable to a dimensionless ratio $x=\nu/\nu_{0}$. Accordingly, all models below are defined in terms of $x$.
@@
Originally proposed to account for spectral breaks (Sieber, 1973), the broken power law model consists of two power-law segments joined at a break frequency $\nu_{b}$:
$S(x)=\begin{cases}bx^{\alpha_{1}},&\textrm{if }\nu\leq\nu_{b},\\ bx^{\alpha_{2}}x_{b}^{\alpha_{1}-\alpha_{2}},&\textrm{otherwise,}\end{cases}$ (4)
where $\alpha_{1}$ and $\alpha_{2}$ are the spectral indices below and above the break, and $x_{b}=\nu_{b}/\nu_{0}$.
The log-parabolic spectrum, used to identify spectral curvature and turn-overs (Kuzmin and Losovsky, 2001; Kijak et al., 2011), is written as
$S(x)=\exp\left(ax^{2}+bx+c\right),$ (5)
@@
Two additional models are derived from a general form that includes an absorption or cut-off term (Sieber, 1973; Malofeev and Malov, 1980):
$S(x)=bx^{\alpha}e^{-\tau(x)}\,.$ (6)
The low-frequency turn-over power law (Jankowski et al., 2018; Rajwade et al., 2016) models the optical depth to produce a low-frequency turn-over, often attributed to thermal free–free absorption (Kijak et al., 2011; Wilson et al., 2013):
$S(x)=bx^{\alpha}\exp\left(\frac{\alpha}{\beta}x_{c}^{-\beta}\right),$ (7)
where $\nu_{c}$ is the turn-over frequency, $x_{c}=\nu_{c}/\nu_{0}$, and $\beta\in[0,2.1]$ controls the smoothness of the transition.
The high-frequency cut-off power law incorporates a high-frequency cut-off by adopting an optical-depth term that produces a linear attenuation with frequency (Kontorovich and Flanchik, 2013). In this case, the optical depth takes the form $\tau(x)=-\ln\!\left(1-x/x_{c}\right)$, leading to
$S(x)=bx^{\alpha}\left(1-\frac{x}{x_{c}}\right),$ (8)
where $\nu_{c}$ is the cut-off frequency and $x_{c}=\nu_{c}/\nu_{0}$. Such high-frequency behaviour is generally interpreted as arising from the pulsar’s interaction with its environment rather than the intrinsic emission mechanism (Kijak et al., 2021; Ardavan, 2024).
Motivated by earlier attempts to combine low- and high-frequency features (Sieber, 1973), the double turn-over spectrum (Swainston et al., 2022) includes both a low-frequency turn-over and a high-frequency cut-off:
$S(x)=bx^{\alpha}\left(1-\frac{x}{x_{c_{1}}}\right)\exp\left(\frac{\alpha}{\beta}x_{c_{2}}^{-\beta}\right),$ (9)
where $\nu_{c_{1}}$ and $\nu_{c_{2}}$ are the cut-off and turn-over frequencies, respectively, with $x_{c_{1}}=\nu_{c_{1}}/\nu_{0}$ and $x_{c_{2}}=\nu_{c_{2}}/\nu_{0}$.
The prior distributions for all model parameters are listed in Table. 1.
@@
Low-frequency turn-over power law $\nu_{c}$ [MHz] $LU(\nu_{\min},\nu_{\max})$
@@
High-frequency cut-off power law $\nu_{c}$ [MHz] $LU(\nu_{\min},1.5\,\nu_{\max})$
@@ III Results
Figure 2: Spectrum of PSR J1809–1917, which was previously classified as a simple power law Jankowski et al. (2018) (lower panel), but our Bayesian analysis identifies the low-frequency turn-over power law as the best model (upper panel), with a $\ln\mathrm{BF}$ of 150 in its favor. The solid orange line and the shaded region display the posterior predictive distribution and its associated uncertainty, with the dark and light colored region showing the 1-$\sigma$ and 3-$\sigma$ credibility interval, respectively. Solid error bars denote the uncertainties reported in the literature, while dashed error bars represent the rescaled uncertainties with inferred $e_{\text{fac}}$ parameters.
@@ III.1 Bayesian model selection
Best-fit model $\ln\mathrm{BF}\geq 5$ (vs. second-best model) $\ln\mathrm{BF}\geq 5$ (vs. simple power law)
@@
As noted in Sec. II.2, the highest-evidence label is not necessarily a unique description. Relying solely on model evidence can be misleading when the differences are small. We therefore compute the log Bayes factor between each curved or broken model and the simple power law. A log Bayes factor of at least 5 is commonly interpreted as decisive support. As shown in Table. 3, 68.8% of pulsars are decisively better fit ($\ln\mathrm{BF}\geq 5$) by a curved or broken model than by a simple power law, indicating that such spectra are common. Even under a more stringent threshold ($\ln\mathrm{BF}\geq 8$), 65.5% of pulsars still favor one of the curved or broken models against the simple power law. Conversely, only 3 of the 121 pulsars that are best fit with a simple power law spectrum has a log Bayes factor over 5 against other spectral models. This suggests that, even when the simple power law yields the highest evidence, its preference is not very strong.
We repeat the analysis using only pulsars with at least ten flux density measurements. With this larger number of data points, the spectral shape is more tightly constrained. Under this stricter requirement, 76.4% (584 out of 764) of categorized pulsars decisively favor ($\ln\text{BF}\geq 5$) curved or broken models over the simple power law, indicating that the prevalence of such spectra is not affected by the small number of measurements. Therefore, we conclude that curved or broken spectra are significantly more prevalent than reported by Jankowski et al. (2018).